OpenStudy (anonymous):

what is the square binomial of x2 + 8x + 16 ?

OpenStudy (anonymous):

you mean, you are asked to factor this?

OpenStudy (anonymous):

did you mean x^2 + 8 x + 16?

OpenStudy (hba):

yes ^^

OpenStudy (anonymous):

(x+4)(x+4)

OpenStudy (anonymous):

yes i ask the factor of a binomial

OpenStudy (anonymous):

to solve for x..obviously x=-4.

OpenStudy (anonymous):

you can use ac factoring.

OpenStudy (anonymous):

note, its in the form ax^2 + bx+c.

OpenStudy (anonymous):

then, identify a and c, and multiply a and c.

OpenStudy (anonymous):

then, identify two numbers where when you multiply you will get ac, and when you add you will get "b" (coefficient of the middle term).

OpenStudy (anonymous):

then, rewrite the middle term using those numbers that you find. :))then, split into two groups, where each group has a gcf, then, factor out the gcf. then factor out the common factor. ta that's it. :))

OpenStudy (anonymous):

let me give you a similar example: x^2+4x+4.so, a = 1, b = 4, and c=4.

OpenStudy (anonymous):

so, ac = 1*4=4, we need two numbers which when we multiply we get 4(ac result), and when we add, we get 4(b , coefficient of "x"). note that, 2*2 = 4, and 2 +2 = 4.

OpenStudy (anonymous):

so, we will rewrite the middle terms using 2 and 2.so, x^2 +4x+4 = x^2+2x+2x+4.

OpenStudy (anonymous):

grouping it, (x^2+2x)+(2x+4), factoring gcf: x(x+2)+2(x+2)

OpenStudy (anonymous):

factoring out the common factor, thus: (x+2)(x+2), or (x+2)^2. that's it! :))